OpenStudy (anonymous):

how to find x in x2+16=8x

OpenStudy (anonymous):

is that 2x or x^2 (x-squared)?

OpenStudy (anonymous):

x squared

OpenStudy (anonymous):

then first you subtract 8x on both sides making it become: x^2 - 8x + 16 = 0 so then you try and simplify this equation so it it becomes: (x-4)(x-4)= 0 then separate them into: (x-4)= 0 and (x-4)=0 then for both of them, add 4 on both sides making x=4. so x = 4. i hope the answer is right and that it helped you :)